# AsEP: Benchmarking Deep Learning Methods for Antibody-specific Epitope Prediction

### v1

@@ 5 WALLE: a graph-based method for epitope prediction
Algorithm MCC Precision Recall AUCROC F1
WALLE 0.210 (0.020) 0.235 (0.018) 0.422 (0.028) 0.635 (0.013) 0.258 (0.018)
@@
Algorithm MCC Precision Recall AUCROC F1
WALLE 0.077 (0.015) 0.143 (0.017) 0.266 (0.025) 0.544 (0.010) 0.145 (0.014)
@@
Experiment results We evaluated each method for both dataset split settings on the test set using the metrics described in Section A.2. Table 1(a) and Table 1(b) summarize the average performance metrics across the test set samples. WALLE generally shows better performance among all metrics except for recall for both dataset splits. We also provided the baseline performance for bipartite link prediction in Table S6.
@@
We also carried out ablation studies (Appendix C) to investigate the impact of different components of WALLE. When we replace the graph convolutional layers with fully connected layers, its performance degenerates considerably, suggesting that the graph convolutional layers contribute to the model’s performance. This is related to the fact that the interaction between a pair of protein structures is dependent on the spatial arrangement of the residues as discussed by Reis et al. (2022). The interface polar bonds, a major source of antibody specificity, tend to shield interface hydrophobic clusters. In addition, the language model embeddings also contribute to the model’s performance, as performance drops when they are replaced by one-hot or BLOSUM62 embeddings. Finally, we also investigated whether the choice of language model affects the model’s performance. We found that the model using AntiBERTy and ESM2 embeddings for antibodies and antigens performed slightly better than the model using ESM2 embeddings for both antibodies and antigens. This suggests that the choice of language model may impact the model’s performance, but a model like ESM2, which is trained on general protein sequences, may contain sufficient information for the epitope prediction task.
@@
While WALLE outperforms other methods in the second dataset split setting, its performance degenerated considerably from the first dataset split setting. This suggests that WALLE is likely biased toward the epitopes in the training set and does not generalize well to unseen epitopes. The performance of the other four methods is not ideal for this task. We believe this would require a more sophisticated model architecture and a more comprehensive dataset to improve the performance of epitope prediction. We invite researchers to examine this issue and bring their insights to improve generalization.
@@
Edge features In terms of structure representation, we only used a simple invariant edge feature, the distance matrix, to capture the neighborhood information of each residue. This topological descriptor already performs better than other methods that use sequence-based features. For future work, more edge features can be incorporated to enrich the graph representation, in addition to the invariant edge features used in this work, such as inter-residue distances and edge types used in GearNet (Zhang et al., 2023), and SE3 equivariant features, such as rotational and orientational relationships between residues as used in abdockgen (Jin et al., 2022).
Antibody types We also plan to extend our work to include other types of antibodies. Current work only looks at conventional antibodies, consisting of heavy- and light-chain variable domains. There are also an increasing number of novel antibodies, for example nanobodies, which are single-variable-domain antibodies derived from camelids. These will be included in future work.
@@ 6 Conclusion
In this work, we proposed a novel benchmarking dataset for the epitope prediction task and clustered the samples by epitopes. We also provided a model, WALLE, which combines protein language models and graph neural networks to leverage their abilities in capturing amino acid contextual and geometric information. We benchmarked WALLE and four other methods, showing that while WALLE outperforms existing methods on both tasks, there remains room for improvement. Nevertheless, we showed the combination of protein language models and graph neural networks improved the performance on the epitope prediction task. We also discussed possible future directions. This work can serve as a starting point for future research in this area.
@@ 7 Acknowledgments
CL was part-funded by a UCL Centre for Digital Innovation Amazon Web Services (AWS) Scholarship; LD was funded by an ISMB MRC iCASE Studentship (MR/R015759/1).
@@ A.1 Related work
Gao et al. 2022 Gao et al. (2022) 258 AbAg complexes
CSM-AB Myung et al. (2021) 472 AbAg complexes
SAGERank Sun et al. (2023) 287 AbAg complexes
Bepipred3.0 Clifford et al. (2022) 582 AbAg complexes
@@
SCEptRe Mahajan et al. (2019) is a related dataset that keeps a weekly updated collection of 3D complexes of epitope and receptor pairs, for example, antibody-antigen, TCR-pMHC, and MHC-ligand complexes derived from the Immune Epitope Database (IEDB). Our approach for clustering antibody-antigen complexes regarding their epitopes is similar to theirs, with the difference in the clustering strategy. We cluster by antigen, then epitope group, and we allow mutated amino acids in the same epitope region because we turn the epitope sites into columns in the multiple sequence alignment. In contrast, SCEptRe clusters by antibody and then compares epitope similarity by epitope conformation using atom-pair distances via PocketMatch Yeturu & Chandra (2008), which is beneficial for comparing the function of various paratopes but is less suitable for our task of predicting epitope residues.
@@
Sequence-based epitope predictor We also tested purely sequence-based epitope prediction tool, for example, Bepipred3.0 Clifford et al. (2022) on our dataset. Bepipred3.0 uses ESM2 model, esm2_t33_650M_UR50D to generate sequence embeddings and was trained on a smaller dataset of 582 antibody-antigen structures and evaluated on 15 antibody-antigen complexes. The authors provided a relatively larger evaluation of linear B-cell epitopes derived from the Immune Epitope Database and reported an AUC-ROC of $0.693$ on the top $10\%$ of the predicted epitopes. We tested Bepipred3.0 on our dataset and found its performance degenerates significantly, as shown in the table below. This is not surprising because linear epitopes are consecutive positions in an antigen sequence, and this task fits better with language model design. Additionally, as pointed out by the authors, approximately $90\%$ of epitopes (B-cell) fall into the conformational category Clifford et al. (2022), which highlights the importance of the present benchmark dataset composed of conformational epitopes derived from filtered antibody-antigen  structures. We believe these results underline the findings in our paper, showing that large language models alone, even if specialized for antibody-antigen interactions, do not encompass all the relevant information needed for epitope prediction.
@@ A.2 Evaluation
In this work, we focus on the epitope prediction task. We evaluate the performance of each method using consistent metrics. Matthew’s Correlation Coefficient (MCC) is highly recommended for binary classification assessments Matthews (1975) and is especially advocated for its ability to provide equal weighting to all four values in the confusion matrix, making it a more informative metric about the classifier’s performance at a given threshold than other metrics Chicco & Jurman (2020, 2023). We encourage the community to adopt MCC for the epitope prediction task as it takes into account true and false positives, as well as true and false negatives, offering a comprehensive measure of the performance. It is considered superior to the AUC-ROC, which is evaluated over all thresholds. For consistency, we also included Precision and Recall from prior studies EpiPred Krawczyk et al. (2014) and PECAN Pittala & Bailey-Kellogg (2020), and we added Area Under the Receiver Operating Characteristic Curve (AUC-ROC) and F1 score, both are typical binary classification metrics. For methods that predict antibody-antigen complex structures, we determine the epitopes using the same distance criterion as aforementioned.
@@ A.3 Steps to build antibody-antigen complex dataset
We sourced our initial dataset from AbDb (version dated September 26, 2022), containing 11,767 antibody files originally collected from the Protein Data Bank (PDB). We collected complexes numbered in Martin scheme Raghavan & Martin (2008) and used AbM CDR definition Martin et al. (1991) to identify CDR residues from the heavy and light chains of antibodies.
@@
To deduplicate complexes, we used MMseqs2 Steinegger & Söding (2017) to cluster the complexes by heavy and light chains in antibodies and antigen sequences. We used the easy-linclust mode with the –cov-mode 0 option to cluster sequences; we used the default setting for coverage of aligned cutoff at 80%; we used different –min-seq-id cutoffs for antibodies and antigens because the antibody framework regions are more conserved than the CDR regions. We cluster heavy and light chains at –min-seq-id cutoff of 100% and 70%, respectively. We retained only one representative file for each unique set of identifiers, leading to a refined dataset of 1,725 files.
@@ A.4 Pipeline to build graph dataset from AbDb
$U$, $V$ and $N$ denote the number of residues in the SEQRES sequence, ATMSEQ sequence and the graph, respectively. thr (at 50 residues) is the cutoff for antigen SEQRES length. We only include antigen sequences with lengths of at least 50 residues. SEQRES and ATMSEQ are two different sequence representations of a protein structure. SEQRES is the sequence of residues in the protein chain as defined in the header section of a PDB file, and it is the complete sequence of the protein chain. ATMSEQ is the sequence of residues in the protein chain as defined in the ATOM section of a PDB file. In other words, it is read from the structure, and any residues in a PDB file are not resolved due to experimental issues that will be missing in the ATMSEQ sequence. Since we are building graph representations using structures, we used ATMSEQ. However, the input to the language models require a complete sequence, therefore we used SEQRES to generate node embeddings, and mapped the node embeddings to the graph nodes. We performed two pairwise sequence alignments to map such embeddings to graph vertices for a protein chain with Clustal Omega Sievers et al. (2011). We first align the SEQRES sequence with the atom sequence (residues collected from the ATOM records in a PDB file) and assign residue embeddings to matched residues. Because we excluded buried residues from the graph, we aligned the sequence formed by the filtered graph vertices with the atom sequence to assign residue embeddings to vertices. ASA is the solvent-accessible surface area of a residue. If a residue has an ASA value of zero, it is considered buried and will be removed from the graph.

### v2

@@ 6 Results and Discussion
Method MCC Precision Recall AUCROC F1
WALLE 0.305 (0.023) 0.308 (0.019) 0.516 (0.028) 0.695 (0.015) 0.357 (0.021)
@@
Method MCC Precision Recall AUCROC F1
WALLE 0.152 (0.019) 0.207 (0.020) 0.299 (0.025) 0.596 (0.012) 0.204 (0.018)
@@
Additionally, we benchmarked AlphaFold2-Multimer (AF2M) version 2.3 on the epitope prediction task due to its growing use in complex structure prediction. A new subset of 76 AsEP complexes, excluded from the AF2M training set, was curated for this purpose. Details on the filtering method, the specific AsEP files in this subset, and AF2M performance results can be found in Section A.9 and Table S5. While AF2M achieves an MCC of $0.262$, its performance could be further improved. Additionally, the average runtime per antibody-antigen pair is $1.66$ hours, which is not optimal for epitope scanning, especially given that all other benchmarked methods here can make predictions within seconds.
Hybrid vs structural only and unsturctured only
@@
We carried out ablation studies (Appendix C) to investigate the impact of different components of WALLE. Specifically, we investigated the combination of GCN with AntiBERTy and ESM-35M for antibody and antigen embeddings. When we replace GCN layers with fully connected layers, the MCC metric decreases by approximately $39.8\%$, suggesting that GNN layers contribute to performance. This is related to the fact that the interaction between a pair of protein structures depends on the spatial arrangement of the residues (see Reis et al. (2022)). The interface polar bonds, a major source of antibody specificity, tend to shield interface hydrophobic clusters. The PLM embeddings also contribute to performance, as performance drops over $62.4\%$ when they are replaced with one-hot or BLOSUM62 embeddings. Finally, we investigated whether the choice of PLM affects the model’s performance. We found that using AntiBERTy and ESM2 embeddings for antibodies and antigens performed slightly better than using ESM2 embeddings for both antibodies and antigens. This suggests that the choice of the protein language model may impact the model’s performance, but a model like ESM2, which is trained on general protein sequences, may contain sufficient information for the epitope prediction task.
WALLE’s model architecture is different from existing approaches in that it combines a graph-based method with pre-trained protein language model embeddings. This integration of structural and unstructured (sequential) information leads to a more comprehensive and information-rich approach to epitope prediction. Unlike WALLE, prior graph-based methods, such as PECAN (Pittala & Bailey-Kellogg, 2020) and EPMP (Vecchio et al., 2021), relied on manually engineered features for node embeddings rather than leveraging pre-trained language models. WALLE’s node embedding update strategy is similar to that of PECAN, where antibody and antigen graphs are updated separately through distinct graph layers. However, EPMP employs GAT layers to jointly update both antibody and antigen embeddings, incorporating a residual connection to add the updated embeddings back to each graph. WALLE also differs in its task formulation and decoder. While PECAN and EPMP primarily focus on node classification tasks, WALLE predicts bipartite link labels and subsequently aggregates them to obtain node labels. Furthermore, WALLE enhances protein language model embeddings with molecular structures, which is different from methods that rely exclusively on sequence-based information, such as ESMBind and ESMFold, or methods that only use structural information, such as EpiPred and MaSIF-site.
@@
Generalization to unseen epitopes While WALLE outperforms other methods in the epitope group dataset split setting, its performance degenerated considerably from the first dataset split setting. This suggests that WALLE is likely biased toward the epitopes in the training set and does not generalize well to unseen epitopes. The performance of the other four methods is not ideal for this task as well. To improve the performance of epitope prediction, we believe this would require a more comprehensive dataset and a more sophisticated model architecture, for example using inductive GNNs (Teru et al., 2020; Zhu et al., 2021; Chen et al., 2022) or pretraining protein language models with active forgetting so that their feature extraction can generalize to unseen epitope, a trick recently proposed to help generalization to unseen languages (Chen et al., 2023). We invite researchers to examine the generalization issue and bring their insights and algorithms to address this issue.
@@
Edge features In terms of structure representation, we only used a simple invariant edge feature, the distance matrix, to capture the neighborhood information of each residue. This topological descriptor already performs better than other methods that use sequence-based features. For future work, more edge features can be incorporated to enrich the graph representation, in addition to the invariant edge features used in this work, such as inter-residue distances and edge types used in GearNet (Zhang et al., 2023), and equivariant features, such as rotational and orientational relationships between residues as used in abdockgen (Jin et al., 2022). The incoporation of edge features will invite testing of advanced graph learning algorithms for multi-relational graphs (Nickel et al., 2011; Bordes et al., 2013; Schlichtkrull et al., 2018; Murphy et al., 2019; Chen et al., 2021).
Dataset We plan to extend our work to include more types of antibodies. Currently, our dataset focuses solely on conventional antibodies composed of heavy-chain and light-chain variable domains. However, there is a growing interest in novel antibody formats, such as nanobodies, which are single-variable-domain antibodies derived from camelids. These will be included in future releases of AsEP. Additionally, given the abundance of general non-antibody-antigen complexes, pre-training on such complexes could be beneficial in providing the downstream epitope prediction models with a foundational understanding of protein interface, as suggested by PECAN (Pittala & Bailey-Kellogg, 2020). Further finetuning on antibody-antigen complexes would then allow the model to capture the unique characteristics of antibody interfaces, such as specific amino acid compositions and interaction types, that distinguish them from general protein interfaces (Kringelum et al., 2013; Ofran et al., 2008). We leave expansions of our dataset to incorporate such diverse data, pretraining over general proteins and finetuning vertically for antibody-antigen complexes as our future work.
@@ 7 Conclusion
In this work, we proposed AsEP, a novel benchmarking dataset for the epitope prediction task and the first dataset to cluster antibody-antigen complexes by epitopes. We also introduced WALLE, a model that combines pretrained protein language models with graph neural networks to capture both amino acid contextual and geometric information. We benchmarked WALLE alongside four other methods, demonstrating that it outperforms existing methods on both tasks, though there is still plenty of room for enhancement, especially on the unseen epitopes. Our results suggest that such integration of both structural modeling and unstructured (sequential) modeling improves epitope prediction performance. We discussed future work that could enhance model performance, including enriching edge features for better structural representation, refining model architecture, exploring transfer learning from general protein complexes, and expanding the dataset to include emerging antibody types. This work serves as our starting point for advancing research in antibody-specific epitope prediction.
@@ 8 Acknowledgments
CL was part-funded by a UCL Centre for Digital Innovation Amazon Web Services (AWS) Scholarship; LD was funded by an ISMB MRC iCASE Studentship (MR/R015759/1). We would also like to thank the NeurIPS reviewers for their valuable feedback and suggestions, which helped us improve the clarity and quality of this work.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes] Limitations are discussed in the paper, although there is no separate section. See Section 5 where the limitations of our graph-based method are discussed, such as its performance degeneration when using epitope-group data splitting, and bias toward seen epitopes. The imbalance of epitope types in the AsEP dataset is also discussed. Furthermore, the limitation to conventional antibodies is made clear (Section 5). Assumptions made, for example, regarding distance cut-offs for epitope definition or prediction thresholds, are either based on established conventions from previous works or optimized for empirically.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] In an impact statement in Appendix D we discuss potential negative impacts of our work. However, this is minimal as it does not directly apply to this field.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] In the abstract section, we provide the link to the GitHub repository with instructions to reproduce our experiments and download the dataset.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] Section A.3: steps to build the dataset; Section 4.3: dataset split; Section A.6: hyperparameters and tunning; Section C.4: details of hyperparameter tuning
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] Table 1 provided standard errors
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] These are provided in Section A.6
    1304923
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] All methods are cited in section Section 2.
    2. (b)
      Did you mention the license of the assets? [Yes] The license of the data set, Creative Commons Attribution 4.0 International, is provided in the data online storage at Zenodo with DOI: 10.5281/zenodo.11495514.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] In the abstract, we provide the link to the GitHub repository with instructions to reproduce our experiments and download the dataset.
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] The original source of the dataset is derived from Protein Data Bank with a CC0 1.0 UNIVERSAL license
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A] The presented dataset does not contain any personally identifiable information or offensive content
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
@@ A.1 Related work
Wang et al. 2022 (Wang et al., 2022) 258 AbAg complexes
CSM-AB (Myung et al., 2021) 472 AbAg complexes
SAGERank (Sun et al., 2023) 287 AbAg complexes
Bepipred3.0 (Clifford et al., 2022) 582 AbAg complexes
In the work by Wang et al. (2022) the dataset from Zhao et al. (2018) was used, which included 257 antibody-antigen complexes, encompassing both VHVL and VHH antibodies. However, the primary focus of this research is on epitope prediction using the antigen as the input. Consequently, antibodies were not included in the predictive models, rendering the dataset unsuitable for antibody-specific epitope prediction.
For CSM-AB, as described by their supplementary information, the dataset contains 472 antibody-antigen structures including 375 Fab, 82 Nanobody and 12 scFv (Myung et al., 2021). These structures were collected from PDB, identified using Chothia annotation as in (Dunbar & Deane, 2015). The authors did not describe the procedure of any further filtering steps. We assumed they included all available structures at that time with available protein-protein binding affinity information from PDBbind since the study aims to predict binding affinity.
The authors of SAGERank formed a dataset composed of 287 antibody-antigen complexes filtered by sequence identity at 95% (Sun et al., 2023) . While the authors did not explicitly mention the antibody types, we infer from the results that these are also Fab antibodies that include both VH and VL domains. The dataset was composed mainly for docking pose ranking output by MegaDock and did not include interface clustering, i.e. epitope grouping.
@@
SCEptRe by  Mahajan et al. (2019) is a related dataset that keeps a weekly updated collection of 3D complexes of epitope and receptor pairs, for example, antibody-antigen, TCR-pMHC, and MHC-ligand complexes derived from the Immune Epitope Database (IEDB). Our approach for clustering antibody-antigen complexes regarding their epitopes is similar to theirs, with the difference in the clustering strategy. We cluster by antigen, then epitope group, and we allow mutated amino acids in the same epitope region because we turn the epitope sites into columns in the multiple sequence alignment. In contrast, SCEptRe clusters by antibody and then compares epitope similarity by epitope conformation using atom-pair distances via PocketMatch (Yeturu & Chandra, 2008), which is beneficial for comparing the function of various paratopes but is less suitable for our task of predicting epitope residues.
@@
Sequence-based epitope predictor We also tested purely sequence-based epitope prediction tool, for example, Bepipred3.0 (Clifford et al., 2022) on our dataset. Bepipred3.0 uses ESM2 model, esm2_t33_650M_UR50D to generate sequence embeddings and was trained on a smaller dataset of 582 antibody-antigen structures and evaluated on 15 antibody-antigen complexes. The authors provided a relatively larger evaluation of linear B-cell epitopes derived from the Immune Epitope Database and reported an AUC-ROC of $0.693$ on the top $10\%$ of the predicted epitopes. We tested Bepipred3.0 on our dataset and found its performance degenerates significantly, as shown in the table below. This is not surprising because linear epitopes are consecutive positions in an antigen sequence, and this task fits better with language model design. Additionally, as pointed out by the authors, approximately $90\%$ of epitopes (B-cell) fall into the conformational category (Clifford et al., 2022), which highlights the importance of the present benchmark dataset composed of conformational epitopes derived from filtered antibody-antigen  structures. We believe these results underline the findings in our paper, showing that large language models alone, even if specialized for antibody-antigen interactions, do not encompass all the relevant information needed for epitope prediction.
@@ A.2 Evaluation
In this work, we focus on the epitope prediction task. We evaluate the performance of each method using consistent metrics. Matthew’s Correlation Coefficient (MCC) is highly recommended for binary classification assessments (Matthews, 1975) and is especially advocated for its ability to provide equal weighting to all four values in the confusion matrix, making it a more informative metric about the classifier’s performance at a given threshold than other metrics (Chicco & Jurman, 2020, 2023). We encourage the community to adopt MCC for the epitope prediction task as it takes into account true and false positives, as well as true and false negatives, offering a comprehensive measure of the performance. It is considered superior to the AUC-ROC, which is evaluated over all thresholds. For consistency, we also included Precision and Recall from prior studies EpiPred (Krawczyk et al., 2014) and PECAN (Pittala & Bailey-Kellogg, 2020), and we added Area Under the Receiver Operating Characteristic Curve (AUC-ROC) and F1 score, both are typical binary classification metrics. For methods that predict antibody-antigen complex structures, we determine the epitopes using the same distance criterion as aforementioned.
@@ A.3 Steps to build antibody-antigen complex dataset
We sourced our initial dataset from AbDb (version dated September 26, 2022), containing 11,767 antibody files originally collected from the Protein Data Bank (PDB). We collected complexes numbered in Martin scheme (Raghavan & Martin, 2008) and used AbM CDR definition (Martin et al., 1991) to identify CDR residues from the heavy and light chains of antibodies.
@@
To deduplicate complexes, we used MMseqs2 (Steinegger & Söding, 2017) to cluster the complexes by heavy and light chains in antibodies and antigen sequences. We used the easy-linclust mode with the –cov-mode 0 option to cluster sequences; we used the default setting for coverage of aligned cutoff at 80%; we used different –min-seq-id cutoffs for antibodies and antigens because the antibody framework regions are more conserved than the CDR regions. We cluster heavy and light chains at –min-seq-id cutoff of 100% and 70%, respectively. We retained only one representative file for each unique set of identifiers, leading to a refined dataset of 1,725 files.
@@ A.4 Pipeline to build graph dataset from AbDb
$U$, $V$ and $N$ denote the number of residues in the SEQRES sequence, ATMSEQ sequence and the graph, respectively. thr (at 50 residues) is the cutoff for antigen SEQRES length. We only include antigen sequences with lengths of at least 50 residues. SEQRES and ATMSEQ are two different sequence representations of a protein structure. SEQRES is the sequence of residues in the protein chain as defined in the header section of a PDB file, and it is the complete sequence of the protein chain. ATMSEQ is the sequence of residues in the protein chain as defined in the ATOM section of a PDB file. In other words, it is read from the structure, and any residues in a PDB file are not resolved due to experimental issues that will be missing in the ATMSEQ sequence. Since we are building graph representations using structures, we used ATMSEQ. However, the input to the language models require a complete sequence, therefore we used SEQRES to generate node embeddings, and mapped the node embeddings to the graph nodes. We performed two pairwise sequence alignments to map such embeddings to graph vertices for a protein chain with Clustal Omega (Sievers et al., 2011). We first align the SEQRES sequence with the atom sequence (residues collected from the ATOM records in a PDB file) and assign residue embeddings to matched residues. Because we excluded buried residues from the graph, we aligned the sequence formed by the filtered graph vertices with the atom sequence to assign residue embeddings to vertices. ASA is the solvent-accessible surface area of a residue. If a residue has an ASA value of zero, it is considered buried and will be removed from the graph.